OpenStudy (anonymous):

solve 2sin2x+root3=0 over the domain [0,2pie]

OpenStudy (anonymous):

(using equivalent angles

OpenStudy (anonymous):

if you want the answer its 2pi/3, 5pi/6, 5pi/3 and 11pi/6

OpenStudy (anonymous):

i understand 2pi/3 and 5pi/6 but dont know how to get the other two solutions

OpenStudy (anonymous):

2sin2x + root3 = 0 2sin2x = (-root3 sin2x = (-root3)/2 this is where it gets a bit tricky. multiply both sides of the equation by sin^-1. sin^-1(sin2x) = sin^-1(-root3)/2 well sin^-1 and sin cancel out so you get 2x on the left. sin^-1 of (-root3)/2 is 4pi/3 and 5pi/3. 2x = 4pi/3 or 5pi/3 divide both by 2 and you get: x = 2pi/6 or 5pi/6. Took me a while to figure it out.

OpenStudy (anonymous):

x = 2pi/3 or 5pi/6. my mistake.

OpenStudy (anonymous):

yeah i got those two solution but apparently theres two more

OpenStudy (anonymous):

read^

OpenStudy (anonymous):

it's a sine wave there's an infinite number of answers. but yeah you mentioned that 0<x<=2pi so yeah 5pi/3 and 11pi/6 would also be answers. and the negatives too I suppose. just graph it on desmos.com

OpenStudy (anonymous):

i dont undestand how 5pi/3 and 11pi/6 are solutions

OpenStudy (anonymous):

you arent allowed a calculator

OpenStudy (anonymous):

your meant to use exact values and equivalent angles. so sin2x = (-root3)/2. Therefore 2x=pi/3

OpenStudy (freckles):

Hey so you agree that we have \[0 \le x \le 2 \pi\] ?

OpenStudy (anonymous):

yess

OpenStudy (freckles):

but inside that sin function we have 2x not x

OpenStudy (anonymous):

yeah

OpenStudy (freckles):

so multiply the inequality by 2 giving you \[0 \le 2x \le 4 \pi\]

OpenStudy (freckles):

let 2x if u so you need to solve sin(u)=-sqrt(3)/2 on the interval [0,4pi]

OpenStudy (freckles):

You will need to find all the solutions that satisfy that on 0 to 2pi then go back around the circle 2pi to 4pi

OpenStudy (freckles):

so you will take the solutions you found in the first interval and then add 2 pi to them

OpenStudy (anonymous):

ok i get now thanks so much

OpenStudy (anonymous):

ok but one more thing

OpenStudy (freckles):

yes?

OpenStudy (anonymous):

why do you multiply the inequality by 2. if its 2x then you solve for u using the normal inequality (0,2pi) and then divide the solutions by 2

OpenStudy (anonymous):

i dont get why you multiply by 2

OpenStudy (freckles):

you solve for u where u is between [0,4pi]

OpenStudy (freckles):

you are asked to solve \[\sin(2x)=\frac{-\sqrt{3}}{2} ,x \in [0, 2\pi] \] If we let u=2x then we have \[\sin(u)=\frac{-\sqrt{3}}{2} , \frac{u}{2} \in [0, 2\pi]\] or rewrite as \[\sin(u)=\frac{-\sqrt{3}}{2} , u \in [0 \cdot 2,2\pi \cdot 2 ] =[0,4\pi]\]

OpenStudy (freckles):

you need to solve that last equation I wrote and find all the solutions that occur in [0,2pi] then to find all the ones in [2pi,4pi] just take the ones you found in the first interval and add 2pi Of course at the end of all of this you will need to replace u with 2x and solve that equation

OpenStudy (anonymous):

ok cheers

OpenStudy (freckles):

You get it?

OpenStudy (freckles):

or still not sure?

OpenStudy (anonymous):

yeah i completely understand how to get the answer. ive just never seen that before sin(2x)=−3√2,x∈[0,2π] If we let u=2x then we have sin(u)=−3√2,u2∈[0,2π] or rewrite as sin(u)=−3√2,u∈[0⋅2,2π⋅2]=[0,4π]

OpenStudy (anonymous):

dw i got it ill just accept that you multiply the domain by 2

OpenStudy (freckles):

Well if u=2x then x=u/2 so that means we have 0<=u/2<=2pi since x was also between 0 and 2pi then to solve the inequality for u you multiply all sides by 2

OpenStudy (freckles):

Let me give an example if you don't mind.

OpenStudy (anonymous):

oh ok i understand now

OpenStudy (anonymous):

thanks

OpenStudy (freckles):

Pretend I was asked to solve sin(3x)=1/2 where 0<=x<=2pi let 3x=u so x=u/3 so we have rewriting this in terms of u sin(u)=1/2 where 0<=u/3<=2pi To solve for u we must multiply all sides by 3 giving the inequality 0<=u<=6pi So looking between [0,2pi] we get u=pi/6 , 5pi/6 then go to [2pi,4pi] (this another full rotation so we also have u=pi/6+2pi , 5pi/6+2pi another full rotation [4pi,6pi] so we have u=pi/6+4pi, 5pi/6+4pi we have 6 solutions in terms of u we will also have 6 solutions in terms of x recall u=3x so we have \[3x=\frac{\pi}{6}, \frac{5\pi}{6},\frac{\pi}{6}+2\pi , \frac{5\pi}{6}+2\pi, \frac{\pi}{6}+4\pi , \frac{5\pi}{6}+4\pi \\ \text{ then we solve for x } \\ x=\frac{\pi}{18} ,\frac{5\pi}{18} ,\frac{\pi}{18} +\frac{2\pi}{3},\frac{5\pi}{18}+\frac{2\pi}{3}, \frac{\pi}{18}+\frac{4\pi}{3}, \frac{5\pi}{18}+\frac{4\pi}{3}\]

OpenStudy (freckles):

So I think the main thing you were having trouble with on the question before is why I multiplied by 2 do you see why I multiply the one inequality in this problem by 3?